OpenStudy (anonymous):

3x+11=y x+y=3 system of linear equations subsitution

OpenStudy (anonymous):

like this? 3x+11x+y=3

OpenStudy (anonymous):

or like this? 3x+11 = ? x+y=3

OpenStudy (anonymous):

3x+11=y

OpenStudy (anonymous):

3x+11=y x+y=3

OpenStudy (anonymous):

substitution means we take one of the two equations, set it equal to 'x' then plug it into the other equation. set the equation x+y=3, equal to x x+y=3 [subtract y from both sides] x+y-y = 3-y x=3-y now substitute this value of 3-y in place of 'x' in the equation 3x+11=y